OpenStudy (anonymous):

If cos x = sin(20 + x)° and 0° < x < 90°, the value of x is ?

OpenStudy (anonymous):

well: 90 - x - 20 = x combine like terms 70-x=x add x to each side 2x = 70 divide by 2 for each side x = 35 to check: cos(35)=.8 sin(35+20)=.8 there we are :) x=35 :)

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

my pleasure :)